# TRACE: Time SeRies PArameter EffiCient FinE-tuning

### v2

@@ 4.4.2 Our solution: Gated Dynamic Simulation Importance Calculation
We propose a method to eliminate bias and precisely compute importance, which is the Two-Stage Gated Dynamic Simulation Importance Calculation (Gated DSIC). This method ensures that the conditions before and after masking remain consistent. The procedure is outlined as follows:
@@
Stage 2: Unbiased Importance Evaluation
@@
    Repeated Sampling: Repeat the random masking operation $K$ times, each time obtaining a different masked set $\tilde{\mathcal{G}}$ and corresponding importance scores $G^{(k)}_{ij}$.
@@
    Average Gradient: Take the average gradient importance across the $K$ samples to obtain an unbiased and stable evaluation:
    $\bar{G}_{ij}=\frac{1}{K}\sum_{k=1}^{K}G^{(k)}_{ij}$ (18)
@@
Based on $\bar{G}_{ij}$, we rank the gates and prune the $K$ gates with the lowest importance (set them to 0), thereby retaining the most efficient LoRA modules.
This method, through multiple random experiments, constructs the post-masking model parameter conditions, ensuring an unbiased calculation of the LoRA Gate importance.
@@ 4.4.3 Workflow
6:        Compute the sensitivity $I^{(m)}(g_{ij})$ for the LoRA gates $g_{ij}\in\mathcal{G}/\tilde{\mathcal{G}}$
@@
8:     Compute the average sensitivity for all LoRA gates: $I(g_{ij})=\frac{1}{M}\sum_{m=1}^{M}I^{(m)}(g_{ij})$ and mask the top $p\%$ ranked gates
@@ 5.1 Datasets and Evaluation Metrics
Datasets. For forecasting and anomaly detection, we used subsets of the MOMENT dataset. Long-term forecasting tasks include ETTh1, ETTh2, ETTm1, ETTm2 [24], Weather, and Exchange [10]. Short-term forecasting uses M4 and M3 from the Monash archive [25], consisting of Yearly (macroeconomic indicators, 5–50 years, horizon=6), Quarterly (sales/economic data, 10–100 quarters, horizon=8), and Monthly (retail/energy/weather data, 50–1000 months, horizon=18). Anomaly detection uses the TSB-UAD benchmark [26], with 1980 univariate time series annotated from 18 public datasets. We selected 9 series from MITDB, ECG, MGAB, and SVDB for experiments.
@@ 5.2 Compared models
[…]
Anomaly Detection: (1) Linear Probing: The model is linear-probed ($\text{MOMENT}_{LP}$) by freezing all parameters except for those in the reconstruction or forecasting head. (2) LoRA Fine-tuning: Based on linear probing, LoRA modules are added to all linear layers for fine-tuning. (3) TRACE: The DISC-LoRA method is applied on top of Linear Probing for anomaly detection.
@@ 5.4.1 Long-horizon Forecasting
Forecasting Horizon $\;\;\mathrm{MOMENT}_{LP}\;\;$ $\mathrm{MOMENT}_{LP+LoRA}$ $\mathrm{MOMENT}_{Proj\;down}$ $\mathrm{MOMENT}_{AdaLoRA}$      TRACE
@@
Static Test: To evaluate the performance of our proposed TRACE model against the state-of-the-art $\mathrm{MOMENT}_{LP}$, we conducted a statistical analysis using the dataset’s actual values and the predictions from both models. Our objective was to assess the accuracy and reliability of these models under various conditions.
To quantify predictive performance, we employed a binary classification framework. A prediction was considered accurate if the absolute error between the predicted and actual values was below a predetermined threshold of 5.0[27]; otherwise, it was deemed inaccurate.
$\begin{bmatrix}158993571&1108\\ 985&4932\end{bmatrix}.$
Computational Cost: Table 5 presents the computational costs of different models on the Electricity dataset, along with experimental results under a prediction horizon of 720 and $\beta$=4. The reported results represent the average of five independent runs.
@@
Forecasting Head linear probing 35.38 387.32 41.70
@@ 5.4.2 Short-horizon Forecasting
For short-horizon forecasting, we adopt the zero-shot setting proposed by Oreshkin [28]. Specifically, we first fine-tune MOMENT on a source dataset with a forecasting head and then assess its performance on a target dataset without any additional fine-tuning. It is consistent with the setup in MOMENT [23].
We evaluated our model’s zero-forecasting performance on the M3 and M4 datasets. We maintained the same settings as in the MOMENT[23], following the dataset splits for training, testing, and forecasting horizons as defined in the M3 and M4 competitions. The evaluation metric used was sMAPE ([29]; [30]). Due to the scarcity of daily, hourly, and weekly frequency data, it is challenging for deep models to fully leverage their zero-shot capabilities in these cases. Table 7 presents the correspondence between source and target datasets. The setup for the Fred dataset follows [23] and [28].
@@ 5.4.3 Anomaly Detection
In the anomaly detection task, we selected subsets of datasets from MITDB, ECG, MGAB, and SVDB. Since the task of anomaly detection involves a relatively small number of parameters in the prediction head, which is not suitable for dimensionality reduction or reconstruction, in this context, TRACE is equivalent to Gated DSIC. Table 8 show that TRACE generally outperforms the linear probing.
@@ 5.4.4 Natural Language Processing Task
We are currently conducting experiments on English datasets in the field of natural language processing. The datasets are sourced from the GLUE benchmark: paraphrase detection (MRPC, QQP), sentiment classification (SST-2), natural language inference (MNLI, RTE, QNLI), and linguistic acceptability (CoLA). We also consider three challenging QA datasets to evaluate the performance of our method on large language models: three commonsense question-answering benchmark tasks—ARC-e, ARC-c [31], and BoolQ [32].
@@
Experimental results on the GLUE benchmark are shown in Table 9, and those on the QA datasets are presented in Table 10. Our method, Gated-DISC, achieves the best or second-best performance with the smallest number of trainable parameters in both settings.
@@ 5.5.1 Impact of the Dimensionality Reduction Factor $\beta$
Figure 2 illustrates the effect of varying the dimensionality reduction factor on MSE under the TRACE method, with a forecast horizon of 192 and the Projection Down reconstruction for different tasks. The analysis shows that the overall fluctuation is minimal, with the maximum fluctuation in a single task not exceeding 1%. The optimal dimensionality reduction factor, $\beta$, typically takes values of 4 or 8. A larger value of $\beta$ reduces the number of trainable parameters in the prediction head, which in turn negatively impacts the model’s forecasting performance. Figure 3 presents the results for a forecast horizon of 720, where similar conclusions can be drawn.
@@
5.5.2 Impact of Different Prediction Heads
@@
Forecasting Horizon $\;\;\mathrm{MOMENT}_{LP}\;\;$ $\mathrm{TRACE}_{Less\;Feature}$ $\mathrm{TRACE}_{Avg\;Pool}$ $\mathrm{TRACE}_{Conv2D}$ $\mathrm{TRACE}_{Proj\;Down}$
@@
The comparison results in Table 11, in nearly all tasks, the TRACE method outperforms Linear Probing or performs equivalently. However, when reconstructing the prediction head with Conv2D, the performance is generally suboptimal, and in most cases, it is inferior to Linear Probing, with a performance gap of within 3.8%. A possible reason for this is that 2D convolution may not be well-suited for extracting features from time series with independent channels, as fluctuations between covariates can introduce interference.
@@ 5.6 Conclusions and future work
This paper proposes an efficient fine-tuning method for time-series foundation models, TRACE: Time Series Parameter Efficient Fine-tuning. It leverages the Gated DSIC method to perform unbiased selection of LoRA modules based on their importance and sensitivity, filtering out unimportant LoRA modules while retaining those that are truly effective in fine-tuning. TRACE demonstrates superior performance in long-term forecasting, short-term forecasting, and anomaly detection tasks.
In long-term forecasting, the prediction head often contains an excessive number of parameters, making it prone to overfitting during fine-tuning and limiting the generalization ability of the foundation model. To address this, we propose various approaches for reconstructing the prediction head, which outperform linear probing in our experiments.
A potential limitation of our approach is that the retained LoRA modules (in terms of proportion and rank) may vary across different downstream tasks and forecasting horizons. One possible solution is to store the trained incremental matrix parameters under different conditions, enabling a plug-and-play mechanism across scenarios. Furthermore, our experiments reveal substantial variability in fine-tuning performance across different tasks and datasets. Exploring the deployment of time-series foundation models in real-world time-series applications and investigating their potential in multimodal learning could be promising directions for future research.
Appendix A Multiple Variants of Reconstructed Prediction Heads and Their Parameter Complexity
@@
Appendix B Computational Resources

### v3

@@ 4.3.2 Our solution: Gated Dynamic Simulation Importance Calculation
As discussed in Section 4.3.1, the standard gradient-based importance evaluation (e.g., as used in AdaLoRA) is biased. This bias arises because the importance score for a LoRA module is calculated under the assumption that all other modules are active. However, during pruning, many other modules are masked (set to zero), which fundamentally alters the model’s context and thus the true contribution of the remaining modules. This inconsistency between the evaluation condition (all modules active) and the deployment condition (many modules masked) leads to an inaccurate and biased importance estimate.
To address this, we propose the Gated Dynamic Simulation Importance Calculation (Gated DSIC) method, which aims to reduce this evaluation bias by simulating the post-pruning context during the importance calculation phase. The core idea is to evaluate a module’s importance not in isolation, but within a dynamically changing coalition of other active and masked modules, thereby better approximating its true contribution in the final pruned model.
Specifically, our method performs a Monte Carlo simulation on the validation set. In each simulation trial, we randomly mask a subset of the LoRA gates. The importance of the remaining unmasked gates is then computed based on their gradients in this specific masked context. By repeating this process many times (M times) and averaging the importance scores across all trials, we obtain a more robust and less biased estimate of each module’s contribution. This procedure ensures that the importance evaluation condition is consistent with the diverse contexts the module will experience after pruning, leading to a more reliable selection of modules to retain.
The superiority of our approach in reducing bias is empirically validated by comparing its importance scores against the gold-standard Shapley Value method, as proposed in [held2022shapley]. Following the protocol in [held2022shapley], we compute the Shapley Values for each LoRA module in a MOMENT-base model fine-tuned on the ETTh1 dataset. We then compare the correlation between these Shapley Values and the importance scores generated by our Gated DSIC method and the gradient-based method used in AdaLoRA. As shown in Table 2 , our method achieves a 63.3% correlation with the Shapley Values, while AdaLoRA’s method shows a significantly lower correlation of -22.1%. This strong alignment with the theoretically sound Shapley Value method demonstrates that Gated DSIC provides a much less biased estimation of module importance compared to conventional gradient-based approaches. This allows our method to achieve performance close to the Shapley-based selection while being computationally feasible (0.25 GPU hours vs. 27 GPU hours for Shapley computation).
Table 2: Correlation of Importance Scores with Shapley Values. We compute the Spearman correlation between the importance scores from different methods and the ground-truth Shapley Values for LoRA modules in a MOMENT-base model fine-tuned on ETTh1.
METHOD CORRELATION WITH SHAPLEY VALUE
Gated DSIC (Ours) 63.3%
AdaLoRA (Gradient-based) -22.1%
To further illustrate this point, we visualize the LoRA module importance scores calculated by the three methods on a MOMENT-base model fine-tuned on ETTh1. As shown in Figure 3, the importance distribution generated by AdaLoRA (left) is highly concentrated, indicating a potential bias towards a small subset of modules. In contrast, the distribution from our Gated DSIC method (middle) exhibits a more diverse and balanced pattern, closely resembling the distribution produced by the Shapley Value method (right), which is considered the gold standard for credit assignment. This visual evidence corroborates our quantitative findings, confirming that Gated DSIC effectively mitigates the evaluation bias inherent in standard gradient-based pruning.
Figure 3: Visualization of LoRA Module Importance Scores.
Left: AdaLoRA (Gradient-based). Middle: Gated DSIC (Ours). Right: Shapley Value. The color intensity represents the magnitude of the importance score, with darker shades indicating higher importance. The x-axis represents different LoRA module types (Query, Key, Value, Output, Gate, Up, Down), and the y-axis represents the Transformer layer number (1-12).
These quantitative and qualitative analyses confirm that Gated DSIC effectively mitigates the evaluation bias inherent in standard gradient-based methods. The detailed procedure of our method is as follows:
@@
Stage 2: Less Biased Importance Evaluation
@@
    Repeated Sampling: Repeat the random masking operation $M$ times, each time obtaining a different masked set $\tilde{\mathcal{G}}$ and corresponding importance scores $I^{(k)}(g_{ij})$.
@@
    Average Gradient: Take the average gradient importance across the $M$ samples to obtain a less biased and stable evaluation:
    $\bar{I}(g_{ij})=\frac{1}{M}\sum_{k=1}^{M}I^{(k)}(g_{ij})$ (12)
Figure 4 visualizes how our method calculates the importance score for each LoRA gate. In each of $M$ independent trials, a random subset of LoRA gates is masked (indicated by grey boxes with ’×’). The gradient magnitude of the unmasked gates is then computed on the validation set. The final importance score for each gate (e.g., 0.31, 0.53) is obtained by averaging its gradient magnitudes across all M trials, as defined in Equation 12. This Monte Carlo simulation ensures that the evaluation context approximates the post-pruning state.
Figure 4: Illustration of the Gated DSIC Importance Scoring Process
@@
Based on $\bar{I}(g_{ij})$, we rank the gates and prune the $p\%$ gates with the lowest importance (set them to 0), thereby retaining the most effective LoRA modules. This process is repeated iteratively over training epochs.
Figure 5 depicts the iterative pruning process over $T$ epochs. At each epoch $t$, we prune an additional $p\%$ of the currently unmasked LoRA gates based on their importance scores from Figure 4. This continues until the total number of pruned gates reaches the predefined budget (e.g., $T\times p\%$), achieving efficient fine-tuning with minimal parameter usage.
This method, through multiple random experiments, constructs model contexts that approximate the post-pruning state, leading to a more reliable and less biased importance estimation for LoRA gates.
Figure 5: Illustration of the Iterative Pruning Schedule
@@ 4.3.3 Workflow
6:   Compute the sensitivity $I^{(k)}(g_{ij})$ for the LoRA gates $g_{ij}\in\mathcal{G}/\tilde{\mathcal{G}}$
@@
8:  Compute the average sensitivity for all LoRA gates: $\bar{I}(g_{ij})=\frac{1}{M}\sum_{k=1}^{M}I^{(k)}(g_{ij})$ and mask the top $p\%$ ranked gates
@@ 5.1 Datasets and Evaluation Metrics
Datasets. For forecasting and anomaly detection, we used subsets of the MOMENT dataset. Long-term forecasting tasks include ETTh1, ETTh2, ETTm1, ETTm2 [etth], Weather, and Exchange [zhou2021informer]. Short-term forecasting uses M4 and M3 from the Monash archive [godahewa2021monash], consisting of Yearly (macroeconomic indicators, 5–50 years, horizon=6), Quarterly (sales/economic data, 10–100 quarters, horizon=8), and Monthly (retail/energy/weather data, 50–1000 months, horizon=18). Anomaly detection uses the TSB-UAD benchmark [tsb_uda], with 1980 univariate time series annotated from 18 public datasets. We selected 9 series from MITDB, ECG, MGAB, and SVDB for experiments.
@@ 5.2 Compared models
Forecasting: (1) Linear Probing: The model is linear-probed ($\text{MOMENT}_{LP}$) by freezing all parameters except for those in the reconstruction or forecasting head. (2) LoRA Fine-tuning: Based on linear probing, LoRA modules are added to all linear layers for fine-tuning. (3) Reconstructed Forecasting Head: The forecasting head is reconstructed using the four methods in sec 4.2: ProjDown,AvgPool,LessFeatureProj\;Down,Avg\;Pool,Less\;Feature and Conv 2D.Conv\;2D. (4) $AdaLoRA$: ProjDownProjDown forecasting head with AdaLoRA method. (5) TRACE: In addition to the reconstructed forecasting head, the Gated DSIC method is incorporated. The main model for TRACE uses the ProjDownProj\;Down method for dimensionality reduction in the forecasting head.
Anomaly Detection: (1) Linear Probing: The model is linear-probed ($\text{MOMENT}_{LP}$) by freezing all parameters except for those in the reconstruction or forecasting head. (2) LoRA Fine-tuning: Based on linear probing, LoRA modules are added to all linear layers for fine-tuning. (3) TRACE: The DSIC-LoRA method is applied on top of Linear Probing for anomaly detection.
@@ 5.4.1 Long-horizon Forecasting
Forecasting
Horizon
$\mathrm{MOMENT}_{LP}\;\;$ $\mathrm{MOMENT}_{LP+LoRA}$ $\mathrm{MOMENT}_{Proj\;down}$ $\mathrm{MOMENT}_{AdaLoRA}$      TRACE
@@
Static Test: We further conducted a statistical significance test (McNemar’s test, p-value=0.0071) on the prediction errors, which confirms that TRACE’s improvements over $\mathrm{MOMENT}_{LP}$ are statistically significant.
The detailed methodology and results of this test are provided in Appendix B
Computational Cost: To address concerns regarding the computational overhead of our Gated DSIC importance evaluation, we provide a detailed breakdown of training and inference costs on the Electricity dataset (prediction horizon = 720, $\beta$ = 4) in Table 6. Our TRACE method requires 942.86 seconds of training time, which is approximately 1.30× that of full LoRA (729.58s) and 1.47× that of AdaLoRA (643.31s). This modest increase in training time is solely attributed to the offline Monte Carlo importance evaluation on the validation set and is deemed acceptable given the consistent performance gains demonstrated in Table 5. Crucially, the inference time of the final pruned TRACE model (38.17s) is nearly identical to that of AdaLoRA (37.95s) and full LoRA (38.36s), confirming that our method imposes no additional latency during deployment. Furthermore, TRACE achieves the lowest backbone parameter count (2.89M) among all PEFT baselines, enhancing its efficiency for resource-constrained scenarios. In summary, the slight offline training overhead is a worthwhile trade-off for a more effective and deployment-ready fine-tuned model.
@@
Forecasting
Head
linear probing 35.38 387.32 41.70
@@ 5.4.2 Short-horizon Forecasting
For short-horizon forecasting, we adopt the zero-shot setting proposed by Oreshkin [shortyuce]. Specifically, we first fine-tune MOMENT on a source dataset with a forecasting head and then assess its performance on a target dataset without any additional fine-tuning. It is consistent with the setup in MOMENT [moment].
We evaluated our model’s zero-forecasting performance on the M3 and M4 datasets. We maintained the same settings as in the MOMENT[moment], following the dataset splits for training, testing, and forecasting horizons as defined in the M3 and M4 competitions. The evaluation metric used was sMAPE ([nbeats]; [wu2022timesnet]). Due to the scarcity of daily, hourly, and weekly frequency data, it is challenging for deep models to fully leverage their zero-shot capabilities in these cases. Table 8 presents the correspondence between source and target datasets. The setup for the Fred dataset follows [moment] and [shortyuce].
@@ 5.4.3 Anomaly Detection
In the anomaly detection task, we selected subsets of datasets from MITDB, ECG, MGAB, and SVDB. Since the task of anomaly detection involves a relatively small number of parameters in the forecasting head, which is not suitable for dimensionality reduction or reconstruction, in this context, TRACE is equivalent to Gated DSIC. Table 9 show that TRACE generally outperforms the linear probing.
@@ 5.4.4 Natural Language Processing Task
We are currently conducting experiments on English datasets in the field of natural language processing. The datasets are sourced from the GLUE benchmark: paraphrase detection (MRPC, QQP), sentiment classification (SST-2), natural language inference (MNLI, RTE, QNLI), and linguistic acceptability (CoLA). We also consider three challenging QA datasets to evaluate the performance of our method on large language models: three commonsense question-answering benchmark tasks—ARC-e, ARC-c [clark2018think], and BoolQ [clark2019boolq].
@@
Experimental results on the GLUE benchmark are shown in Table 10, and those on the QA datasets are presented in Table 11. Our method, Gated-DSIC, achieves the best or second-best performance with the smallest number of trainable parameters in both settings.
@@ 5.5.1 Impact of the Dimensionality Reduction Factor $\beta$
Figure 6 illustrates the effect of varying the dimensionality reduction factor on MSE under the TRACE method, with a forecast horizon of 192 and the Projection Down reconstruction for different tasks. The analysis shows that the overall fluctuation is minimal, with the maximum fluctuation in a single task not exceeding 1%. The optimal dimensionality reduction factor, $\beta$, typically takes values of 4 or 8. A larger value of $\beta$ reduces the number of trainable parameters in the forecasting head, which in turn negatively impacts the model’s forecasting performance. Figure 7 presents the results for a forecast horizon of 720, where similar conclusions can be drawn.
@@
5.5.2 Impact of Different Forecasting Heads
@@
Forecasting
Horizon
$\mathrm{MOMENT}_{LP}\;\;$ $\mathrm{TRACE}_{Less\;Feature}$ $\mathrm{TRACE}_{Avg\;Pool}$ $\mathrm{TRACE}_{Conv2D}$ $\mathrm{TRACE}_{Proj\;Down}$
@@
The comparison results in Table 12, in nearly all tasks, the TRACE method outperforms Linear Probing or performs equivalently. However, when reconstructing the forecasting head with Conv2D, the performance is generally suboptimal, and in most cases, it is inferior to Linear Probing, with a performance gap of within 3.8%. A possible reason for this is that 2D convolution may not be well-suited for extracting features from time series with independent channels, as fluctuations between covariates can introduce interference.
@@ 5.6 Conclusions and future work
In this work, we presented TRACE, a parameter-efficient fine-tuning framework tailored for time series foundation models. Our approach addresses two critical bottlenecks: the parameter inefficiency of standard forecasting heads and the biased importance evaluation in existing PEFT methods. Through extensive experiments, we demonstrated that TRACE’s reconstructed heads and the Gated DSIC mechanism consistently yield superior performance across a range of tasks.
A current limitation is that the optimal set of LoRA modules may vary across tasks. For future work, we plan to investigate task-agnostic or meta-learning-based strategies to identify a universal set of adaptable modules. Furthermore, we aim to explore the integration of TRACE with multimodal time series (e.g., combining sensor data with textual logs) to unlock new applications in real-world scenarios.
Appendix A Multiple Variants of Reconstructed Forecasting Heads and Their Parameter Complexity
@@
Appendix B Statistical Validation of Performance
To evaluate the performance of our proposed TRACE model against the state-of-the-art $\mathrm{MOMENT}_{LP}$, we conducted a statistical analysis using the dataset’s actual values and the predictions from both models. Our objective was to assess the accuracy and reliability of these models under various conditions.
To quantify predictive performance, we employed a binary classification framework. A prediction was considered accurate if the absolute error between the predicted and actual values was below a predetermined threshold of 5.0[jiang2025tcm]; otherwise, it was deemed inaccurate.
$\begin{bmatrix}158993571&1108\\ 985&4932\end{bmatrix}$.
Appendix C Computational Resources